OpenStudy (anonymous):

A plane flies 2,000 miles from Chicago to Los Angeles wit ha 50 mph tail wind in 3.25 hours. Returning against that same wind, it takes 4 hours. What is the speed of the plane if there is no wind?

OpenStudy (lgbasallote):

downstream speed = 2,000/3.25 downstream speed = speed of plane + tail wind upstream speed = 2,000/4 upstream speed = speed of plane - tail wind does that help?

OpenStudy (anonymous):

Since your the only person i know that understands it, then yes it helps a lot man!

OpenStudy (lgbasallote):

lol